OpenStudy (anonymous):

Convert 376 cg to kg using dimensional analysis. (Hint: 100 cg = 1 g; 1,000 g = 1 kg) A. 3.76 × 10^–3 kg B. 3.76 × 10^–1 kg C. 3.76 × 10^0 kg D. 3.76 × 10^4 kg E. 3.76 × 10^7 kg

OpenStudy (anonymous):

Can anyone please help?

OpenStudy (mayankdevnani):

what you think what is the answer?

OpenStudy (mayankdevnani):

@RH

OpenStudy (anonymous):

I am not sure what should I convert first

OpenStudy (mayankdevnani):

first convert 376 cg into grams....so you divide 376 by 100 to get grams.. \[\huge \frac{376}{100}=???\]

OpenStudy (anonymous):

3.76

OpenStudy (mayankdevnani):

correct

OpenStudy (anonymous):

now from grams to kg right?

OpenStudy (mayankdevnani):

right

OpenStudy (mayankdevnani):

so what is the answer?

OpenStudy (mayankdevnani):

@RH

OpenStudy (anonymous):

3.76 / 1 kg right (I am not sure...)

OpenStudy (mayankdevnani):

no!! just divide 3.76 by 1000 to get kilograms: \[\huge \frac{3.76}{1000}=0.00376=3.76 \times 10^{-3}\]

OpenStudy (mayankdevnani):

so option a is correct answer

OpenStudy (mayankdevnani):

understood or not? @RH

OpenStudy (anonymous):

Yes yes I understand! Thanks a lot :)

OpenStudy (mayankdevnani):

welcome:)

OpenStudy (anonymous):

Can you help me with one more? I don't know how to start with it ... :(